OpenStudy (anonymous):

i dont understand this Solve the equation S/R + 3 = T/R for R in terms of S and T.

OpenStudy (anonymous):

@jbug55: \[\frac{ S }{ R } + 3 = \frac{ T }{ R }\] In the above eqn. we have to make R the subject of the formula expressed in terms of S and T SO: \[\frac{ S + 3R }{ R } = \frac{ T }{ R}\] Now Transposing: \[S + 3R = \frac{ T }{ R } \times R\] \[S + 3R = T\] \[3R = T -S\] \[R = \frac{ T -S }{ 3 }\]

OpenStudy (anonymous):

I have to point out one tiny constraint to @Adjax's solution - given the original problem R cannot equal zero. Which means in the final solution we have to add the constraint. \[T \neq S\]